# HugAgent: Evaluating LLMs in Simulating Individual-Level Human Reasoning on Open-Ended Tasks

### v2

@@ Appendix C HugAgent Benchmark Details
Model Context Belief-State Inference (Accuracy $\uparrow$) Belief Dynamic Update (Accuracy $\uparrow$) Belief Dynamic Update (MAE $\downarrow$) Health Surv. Zone Avg Health Surv. Zone Avg Health Surv. Zone Avg GPT-4o 5 QAs 68.89±0.51 73.93±1.07 72.22±0.00 71.68 61.48±0.35 67.80±0.45 65.82±0.39 64.22 1.26±0.01 1.19±0.01 1.24±0.00 1.23 10 QAs 67.59±0.17 72.30±0.37 72.06±0.66 70.65 62.29±0.33 64.56±0.39 65.82±0.39 64.22 1.28±0.00 1.26±0.01 1.21±0.01 1.25 20+ QAs 71.11±0.77 75.25±0.90 77.62±0.35 74.66 61.57±0.19 63.74±0.58 64.04±0.20 63.12 1.27±0.01 1.35±0.01 1.26±0.01 1.29 Gemini 2.0 Flash 5 QAs 61.30±0.41 65.25±0.45 65.08±0.00 63.88 59.70±0.09 65.71±0.30 62.25±0.21 62.55 1.36±0.00 1.23±0.00 1.34±0.00 1.31 10 QAs 60.19±0.65 64.75±0.00 71.27±0.35 65.40 58.22±0.28 62.58±0.23 65.20±0.16 61.97 1.37±0.00 1.30±0.01 1.26±0.00 1.31 20+ QAs 63.89±0.00 68.20±0.37 77.78±0.00 69.96 56.23±0.12 62.25±0.37 63.16±0.10 60.55 1.39±0.00 1.33±0.00 1.33±0.00 1.35 Qwen2.5-32B-instr 5 QAs 68.33±1.21 73.11±2.68 65.08±0.56 68.84 59.45±0.55 62.53±0.57 63.38±0.69 61.79 1.36±0.00 1.31±0.01 1.27±0.01 1.31 10 QAs 75.37±4.47 75.90±1.80 67.78±1.06 73.02 58.52±0.74 60.66±0.60 63.02±0.47 60.73 1.41±0.01 1.36±0.01 1.27±0.01 1.35 20+ QAs 77.96±0.41 76.56±0.45 76.98±0.79 77.17 56.82±0.41 57.80±0.25 62.25±0.16 58.96 1.43±0.00 1.43±0.01 1.33±0.01 1.40
Table 10: Human belief-state inference and belief dynamic update performance across question lengths (5Q, 10Q, 20Q) for Gemini 2.0 Flash, Qwen2.5-32B-instr, and GPT-4o. Reported as mean $\pm$ std over 5 runs and across domains (healthcare, surveillance, zoning). Best in each column (across all models) is bolded.
Task Model Health $\rightarrow$ Surv Surv $\rightarrow$ Zone Zone $\rightarrow$ Health Avg Belief-State Inference (Accuracy $\uparrow$) Gemini 2.0 Flash 29.26±0.83 31.31±0.69 27.30±1.91 29.29 Qwen2.5-32B-instr 53.52±1.52 58.69±1.10 55.08±1.20 55.76 GPT-4o 58.52±0.77 65.41±1.58 51.75±1.03 58.56 Belief Dynamic Update (Accuracy $\uparrow$) Gemini 2.0 Flash 45.34±0.21 40.77±0.15 46.84±0.28 44.31 Qwen2.5-32B-instr 44.41±0.98 24.84±0.60 30.15±0.35 33.13 GPT-4o 49.49±0.87 42.47±0.50 41.96±0.30 44.64 Belief Dynamic Update (MAE $\downarrow$) Gemini 2.0 Flash 1.70±0.00 1.70±0.00 1.74±0.00 1.71 Qwen2.5-32B-instr 1.69±0.01 1.94±0.01 2.05±0.01 1.89 GPT-4o 1.55±0.01 1.60±0.00 1.71±0.00 1.62
Table 11: Human Cross_Topic test: models are trained in one topic domain and evaluated on another. Reported as mean $\pm$ std over 5 runs and across domains (healthcare, surveillance, zoning). All accuracy metrics are shown as percentages.
Task Model Belief-State Inference (Accuracy $\uparrow$) Belief Dynamic Update (Accuracy $\uparrow$) Belief Dynamic Update (MAE $\downarrow$) Health Surv. Zone Avg Health Surv. Zone Avg Health Surv. Zone Avg Swap-cross_person Gemini 2.0 Flash 52.78±0.65 52.79±1.49 47.46±0.66 51.01 34.19±0.12 40.71±0.36 39.35±0.16 38.08 2.13±0.01 1.85±0.01 2.05±0.00 2.01 Qwen2.5-32B-instr 58.15±0.77 64.75±0.00 58.41±0.90 60.44 33.77±0.73 40.77±0.63 38.30±3.63 37.61 2.19±0.01 1.87±0.01 2.04±0.06 2.03 GPT-4o 50.93±7.29 61.31±0.37 51.11±0.71 54.45 37.37±0.19 43.13±0.75 37.38±0.28 39.29 2.02±0.01 1.80±0.01 1.98±0.01 1.93
@@
Table 12: Human Cross_Person test: models trained on one participant and evaluated on another. Reported as mean $\pm$ std over 5 runs and across domains (healthcare, surveillance, zoning).
@@
Model Context Belief-State Inference (Accuracy $\uparrow$) Belief Dynamic Update (Accuracy $\uparrow$) Belief Dynamic Update (MAE $\downarrow$) Health Surv. Zone Avg Health Surv. Zone Avg Health Surv. Zone Avg Gemini 2.0 Flash 5 QAs 37.97±0.00 63.48±0.60 57.38±0.00 52.94 42.75±0.18 63.64±0.14 54.09±0.18 53.49 1.82±0.00 1.18±0.00 1.54±0.00 1.51 10 QAs 43.04±0.00 61.74±0.49 62.13±0.90 55.64 41.79±0.25 63.02±0.14 55.73±0.14 53.51 1.83±0.00 1.19±0.00 1.52±0.00 1.51 20+ QAs 51.90±0.00 63.48±1.24 63.61±0.45 59.66 45.28±0.20 62.96±0.00 55.93±0.22 54.72 1.77±0.00 1.18±0.00 1.49±0.00 1.48 Qwen2.5-32B-instr 5 QAs 47.34±2.12 67.61±1.19 66.56±0.37 60.50 45.37±0.57 63.70±0.71 53.00±0.54 54.02 1.84±0.01 1.18±0.01 1.56±0.01 1.53 10 QAs 59.24±1.39 68.70±0.91 70.82±0.73 66.25 45.37±0.56 65.00±0.17 55.88±0.89 55.42 1.83±0.01 1.18±0.01 1.52±0.01 1.51 20+ QAs 64.30±0.57 68.70±0.49 71.64±0.93 68.21 45.37±0.39 65.00±0.17 55.93±0.54 55.37 1.80±0.00 1.18±0.00 1.50±0.01 1.49
@@
Table 13: Synthetic belief-state inference and belief dynamic update performance across question lengths (5Q, 10Q, 20Q) for Gemini 2.0 Flash and Qwen2.5-32B-instr. Reported as mean $\pm$ std over 5 runs and across domains (healthcare, surveillance, zoning). Best in each column (across all models) is bolded.
@@
Task Model Health $\rightarrow$ Surv Surv $\rightarrow$ Zone Zone $\rightarrow$ Health Avg Belief-State Inference (Accuracy $\uparrow$) Gemini 2.0 Flash 17.97±0.0226 38.91±0.0091 20.49±0.0130 25.79 Qwen2.5-32B-instr 22.78±0.0155 48.26±0.0197 45.57±0.0073 38.87 Belief Dynamic Update (Accuracy $\uparrow$) Gemini 2.0 Flash 36.03±0.0022 41.73±0.0051 52.06±0.0032 43.27 Qwen2.5-32B-instr 34.28±0.0056 23.40±0.0034 26.65±0.0089 28.11 Belief Dynamic Update (MAE $\downarrow$) Gemini 2.0 Flash 1.97±0.00 1.70±0.01 1.62±0.01 1.76 Qwen2.5-32B-instr 2.16±0.01 1.88±0.01 2.12±0.01 2.05
@@
Table 14: Synthetic Cross_Topic test: models are trained in one topic domain and evaluated on another. Reported as mean $\pm$ std over 5 runs and across domains (healthcare, surveillance, zoning).
@@
Task Model Belief-State Inference (Accuracy $\uparrow$) Belief Dynamic Update (Accuracy $\uparrow$) Belief Dynamic Update (MAE $\downarrow$) Health Surv. Zone Avg Health Surv. Zone Avg Health Surv. Zone Avg Swap-cross_person Gemini 2.0 Flash 30.89±0.69 47.83±0.00 33.93±0.45 37.55 30.13±0.15 50.93±0.00 35.53±0.11 38.86 2.16±0.00 1.52±0.00 2.06±0.00 1.91 Qwen2.5-32B-instr 42.28±1.13 55.00±0.60 34.75±0.45 44.01 32.10±0.53 52.78±0.22 35.04±0.32 39.97 2.14±0.01 1.49±0.00 2.06±0.00 1.90
@@
Table 15: Synthetic Cross_Person test: models trained on one participant and evaluated on another. Reported as mean $\pm$ std over 5 runs and across domains (healthcare, surveillance, zoning).
@@ Appendix H Universal Healthcare Questionnaire
The survey design draws on the four-stage cognitive model of survey response (Tourangeau et al., 2000):
@@ Appendix I Surveillance Camera Questionnaire
The survey design draws on the four-stage cognitive model of survey response (Tourangeau et al., 2000):
@@
Our benchmark targets reasoning as a dynamic and individualized process, rather than static prediction. We therefore adopt an open-ended elicitation principle: instead of pre-defining fixed question banks, HugAgent uses a single guiding question to initiate a semi-structured conversation. All follow-up questions are generated adaptively within the same dialogue, grounded in the participant’s own responses. This design enables deep, conversational reasoning to unfold while minimizing artificial scaffolding from the chatbot itself. We do not claim fully open-world coverage; rather, we emphasize open-domain extensibility: by simply swapping the guiding question, the benchmark can be ported to new domains while maintaining consistency in evaluation. Such minimal-interaction protocols align with prior work showing that lightweight conversational scaffolds preserve ecological validity in human reasoning studies (Ericsson & Simon, 1980; Clark, 1996; Sap et al., 2019b; Driess et al., 2023). This principle directly motivates the guiding-question chatbot protocol we describe in Appendix D.
@@
To evaluate whether models capture not only what individuals believe but also how their beliefs evolve, we operationalize reasoning through two proxy tasks: belief-state inference (recovering stance and factor polarity from context) and belief-dynamics update (predicting stance shifts and reweighting under new evidence). These tasks follow the tradition of modeling belief revision as a tractable proxy for underlying cognitive processes (Gopnik & Schulz, 2004; Sloman, 2009). While other proxies could be envisioned, these two are the most direct operationalizations of individual reasoning trajectories, balancing interpretability and task difficulty. This motivates our benchmark’s two-task structure, detailed in Appendix D.
@@
Human data provide ecological validity: rich, idiosyncratic reasoning paths embedded in natural language. Synthetic agents, by contrast, provide controllability and scale: fully specified stance profiles and deterministic update rules allow stress-testing model adaptation under known ground truth. Together, the two tracks are complementary: humans as ecological baselines, synthetics as controlled stress tests. This mirrors dual-track designs in cognitive science and simulation benchmarks, where naturalistic and synthetic data jointly enhance validity and reproducibility (Lake et al., 2017; Battaglia et al., 2018). Synthetic agents are not intended to replace human data but to serve as a complementary axis of evaluation. This dual-track design is what anchors HugAgent between ecological realism and controlled generalization tests.
@@
Instead, we adopt test–retest reliability as the human ceiling. Here, the same participant is re-sampled or re-interviewed, and the consistency of their own responses provides a direct measure of reliability. This practice is well established in psychology and survey research, and has been adopted in recent large-scale reasoning datasets facing similar challenges (Park et al., 2023; Toubia et al., 2025). Compared to annotator baselines, test–retest reliability offers a more precise and ecologically valid upper bound for model performance, aligned with the benchmark’s goal of capturing intra-individual reasoning fidelity. This principle defines how we report the human ceiling in HugAgent.

### v3

@@ Appendix C HugAgent Benchmark Details
Model Belief Dynamics Update
Change Detection Acc. ($\uparrow$) Direction Inference Acc. ($\uparrow$) Dir. Acc. ($\uparrow$)
[…]
OpenAI Models
GPT-4o 49.36±1.59 65.83±3.12 57.14±3.01 88.89±0.00 91.49±0.53 98.33±3.33 77.03±0.48 83.79±1.31 85.98±1.71
GPT-5-mini 46.81±4.66 65.83±6.12 60.95±3.56 82.95±2.56 90.67±5.33 82.05±9.48 72.11±2.84 83.22±3.06 75.72±7.17
o3-mini 50.64±4.54 55.83±4.25 59.05±5.71 76.43±4.84 79.58±3.64 78.59±12.06 68.69±3.33 72.45±3.10 72.72±8.67
Other Closed-Source Models
Gemini 2.0 Flash 38.30±0.00 70.83±0.00 62.86±1.90 100.00±0.00 83.33±0.00 100.00±0.00 81.49±0.00 79.58±0.00 88.86±0.57
DeepSeek-R1-0528 54.47±2.17 64.67±5.20 52.38±3.01 83.62±2.10 84.42±3.75 100.00±0.00 74.87±1.86 78.50±1.50 85.71±0.90
Qwen-plus_2025-07-28 34.04±2.33 63.01±1.94 66.67±0.00 85.24±0.95 83.03±0.61 92.31±0.00 69.88±0.97 77.02±0.89 84.62±0.00
Qwen-max_2024-10-15 34.89±1.04 63.33±1.67 66.67±0.00 85.71±0.00 82.05±2.56 92.31±0.00 70.47±0.31 76.44±1.29 84.62±0.00
Open-Source Models
LLaMA_3.3_70B 68.09±3.01 54.17±0.00 53.33±1.90 80.00±0.00 85.71±0.00 100.00±0.00 76.43±0.90 76.25±0.00 86.00±0.57
Qwen2.5-32B-instr. 33.19±1.04 62.50±0.00 66.67±3.01 84.29±1.17 83.33±0.00 92.29±0.38 68.96±1.05 77.08±0.00 84.60±1.17
Qwen2.5-7B-instr 33.62±1.59 62.50±0.00 65.63±1.40 84.29±1.17 83.33±0.00 93.57±3.26 69.09±1.03 77.08±0.00 85.19±2.30
Memory-Augmented Baselines
RAG [^1] RAG (lewis2020retrieval), RAG-FC = RAG with Full Context (lewis2020retrieval). 42.55±1.90 55.83±5.00 66.67±0.00 66.67±0.00 86.51±3.85 85.71±0.00 59.43±0.57 77.30±1.94 80.00±0.00
RAG-FC [^1] RAG (lewis2020retrieval), RAG-FC = RAG with Full Context (lewis2020retrieval). 45.11±2.08 62.43±6.00 60.95±1.90 80.00±0.00 90.91±0.00 86.03±2.82 69.53±0.63 82.36±1.80 78.50±1.40
Generative Agents [^2] Generative Agents (park2023thousandpeople). 36.17±1.90 61.06±2.27 60.95±1.90 81.00±9.70 83.64±3.64 93.85±3.08 67.55±6.41 76.86±2.20 83.98±1.58
[…]
@@
Table 11: Cross-domain swap test: models are trained with QA context from one domain and evaluated on another domain for the same participant. We report performance for Healthcare $\rightarrow$ Surveillance, Surveillance $\rightarrow$ Zoning, Zoning $\rightarrow$ Healthcare, and their average. Reported as mean $\pm$ std over 5 runs.
[…]
Table 12: Question masking / length scaling for both Belief State Inference and Belief Dynamics Update. Reported as mean $\pm$ std over 5 runs. Best results per column are highlighted in bold.
[…]
Table 13: Comparison of No-Context (population prior) and Full-Context (with individual transcripts) settings. Reported as mean $\pm$ std over 5 runs.
Task Model Belief State Inference (Accuracy (% $\uparrow$)) Belief Dynamics Update (Accuracy (% $\uparrow$)) Belief Dynamics Update (MAE $\downarrow$) Health Surv. Zone Avg Health Surv. Zone Avg Health Surv. Zone Avg Swap-cross_person Gemini 2.0 Flash 52.78±0.65 52.79±1.49 47.46±0.66 51.01±2.76 34.19±0.12 40.71±0.36 39.35±0.16 38.08±2.91 2.13±0.01 1.85±0.01 2.05±0.00 2.01±0.12 Qwen2.5-32B-instr. 58.15±0.77 64.75±0.00 58.41±0.90 60.44±3.22 33.77±0.73 40.77±0.63 39.93±0.46 38.16±3.28 2.19±0.01 1.87±0.01 2.01±0.00 2.02±0.13 GPT-4o 50.93±7.29 61.31±0.37 51.11±0.71 54.45±6.37 37.37±0.19 43.13±0.75 37.38±0.28 39.30±2.84 2.02±0.01 1.80±0.01 1.98±0.01 1.93±0.10
Table 14: Human Cross_Person test: models trained on one participant and evaluated on another. Reported as mean $\pm$ std over 5 runs.
@@
Model Context Belief State Inference (Accuracy (% $\uparrow$)) Belief Dynamics Update (Accuracy (% $\uparrow$)) Belief Dynamics Update (MAE $\downarrow$) Health Surv. Zone Avg Health Surv. Zone Avg Health Surv. Zone Avg GPT-4o 5 QAs 56.96±1.27 73.48±0.60 62.30±1.00 64.25±7.18 53.23±0.71 64.20±0.00 56.28±0.27 57.90±4.80 1.68±0.00 1.18±0.01 1.43±0.01 1.43±0.21 10 QAs 62.28±1.06 74.13±0.49 66.89±0.45 67.76±5.09 52.18±0.44 65.12±0.44 55.53±1.78 57.61±5.76 1.73±0.01 1.17±0.01 1.44±0.00 1.45±0.24 20+ QAs 60.76±0.00 75.00±0.77 68.03±0.58 67.93±6.04 52.71±0.29 65.00±0.35 58.11±0.48 58.61±5.22 1.72±0.00 1.18±0.01 1.41±0.01 1.44±0.23 Gemini 2.0 Flash 5 QAs 37.97±0.00 63.48±0.60 57.38±0.00 52.94±11.26 42.75±0.18 63.64±0.14 54.09±0.18 53.49±8.84 1.82±0.00 1.18±0.00 1.54±0.00 1.51±0.27 10 QAs 43.04±0.00 61.74±0.49 62.13±0.90 55.64±9.24 41.79±0.25 63.02±0.14 55.73±0.14 53.51±9.12 1.83±0.00 1.19±0.00 1.52±0.00 1.51±0.27 20+ QAs 51.90±0.00 63.48±1.24 63.61±0.45 59.66±5.73 45.28±0.20 62.96±0.00 55.93±0.22 54.73±7.52 1.77±0.00 1.18±0.00 1.49±0.00 1.48±0.25 Qwen2.5-32B-instr. 5 QAs 47.34±2.12 67.61±1.19 66.56±0.37 60.50±9.73 45.37±0.57 63.70±0.71 53.00±0.54 54.02±7.80 1.84±0.01 1.18±0.01 1.56±0.01 1.53±0.28 10 QAs 59.24±1.39 68.70±0.91 70.82±0.73 66.25±5.30 44.76±0.41 64.75±0.14 54.74±0.62 54.75±8.46 1.83±0.01 1.18±0.01 1.52±0.01 1.51±0.28 20+ QAs 64.30±0.57 68.70±0.49 71.64±0.93 68.21±3.18 45.24±0.39 65.00±0.17 55.88±0.54 55.37±8.37 1.80±0.00 1.18±0.00 1.50±0.01 1.49±0.26
@@
Table 15: Synthetic belief state inference and belief dynamics update performance across question lengths (5Q, 10Q, 20Q) for GPT-4o, Gemini 2.0 Flash and Qwen2.5-32B-instr.. Reported as mean $\pm$ std over 5 runs. Best in each column (across all models) is bolded.
@@
Task Model Health $\rightarrow$ Surv Surv $\rightarrow$ Zone Zone $\rightarrow$ Health Avg. Belief State Inference (Accuracy (% $\uparrow$)) GPT-4o 18.99±0.90 50.22±0.49 39.67±0.73 36.29±13.44 Gemini 2.0 Flash 17.97±2.26 38.91±0.91 20.49±1.30 25.79±9.77 Qwen2.5-32B-instr. 22.78±1.55 48.26±1.97 45.57±0.73 38.87±11.91 Belief Dynamics Update (Accuracy (% $\uparrow$)) GPT-4o 43.62±0.71 55.43±0.64 35.58±0.28 44.88±8.45 Gemini 2.0 Flash 36.03±0.22 41.73±0.51 52.06±0.32 43.27±6.88 Qwen2.5-32B-instr. 34.28±0.56 23.40±0.34 26.65±0.89 28.11±4.76 Belief Dynamics Update (MAE $\downarrow$) GPT-4o 1.64±0.01 1.45±0.01 1.73±0.00 1.61±0.12 Gemini 2.0 Flash 1.97±0.00 1.70±0.01 1.62±0.01 1.76±0.15 Qwen2.5-32B-instr. 2.16±0.01 1.88±0.01 2.12±0.01 2.05±0.13
@@
Table 16: Synthetic Cross_Domain test: models are trained in one topic domain and evaluated on another. Reported as mean $\pm$ std over 5 runs.
@@
Task Model Belief State Inference (Accuracy (% $\uparrow$)) Belief Dynamics Update (Accuracy (% $\uparrow$)) Belief Dynamics Update (MAE ($\downarrow$) Health Surv. Zone Avg Health Surv. Zone Avg Health Surv. Zone Avg Swap-cross_person GPT-4o 41.52±1.06 54.35±0.00 33.11±0.73 42.99±9.06 40.17±0.22 52.90±0.35 36.72±0.50 43.27±7.21 2.04±0.00 1.54±0.00 2.00±0.00 1.86±0.24 Gemini 2.0 Flash 30.89±0.69 47.83±0.00 33.93±0.45 37.55±7.64 30.13±0.15 50.93±0.00 35.53±0.11 38.86±9.12 2.16±0.00 1.52±0.00 2.06±0.00 1.91±0.29 Qwen2.5-32B-instr. 42.28±1.13 55.00±0.60 34.75±0.45 44.01±8.68 32.10±0.53 52.78±0.22 35.04±0.32 39.97±9.46 2.14±0.01 1.49±0.00 2.06±0.00 1.90±0.30
@@
Table 17: Synthetic Cross_Person test: models trained on one participant and evaluated on another. Reported as mean $\pm$ std over 5 runs.
@@ Appendix H Universal Healthcare Questionnaire
The survey design draws on the four-stage cognitive model of survey response (tourangeau2000psychology):
@@ Appendix I Surveillance Camera Questionnaire
The survey design draws on the four-stage cognitive model of survey response (tourangeau2000psychology):
@@
Our benchmark targets reasoning as a dynamic and individualized process, rather than static prediction. We therefore adopt an open-ended elicitation principle: instead of pre-defining fixed question banks, HugAgent uses a single guiding question to initiate a semi-structured conversation. All follow-up questions are generated adaptively within the same dialogue, grounded in the participant’s own responses. This design enables deep, conversational reasoning to unfold while minimizing artificial scaffolding from the chatbot itself. We do not claim fully open-world coverage; rather, we emphasize open-domain extensibility: by simply swapping the guiding question, the benchmark can be ported to new domains while maintaining consistency in evaluation. Such minimal-interaction protocols align with prior work showing that lightweight conversational scaffolds preserve ecological validity in human reasoning studies (thinkaloud1980; clark1996using; sap2019socialiqa; driess2023palm). This principle directly motivates the guiding-question chatbot protocol we describe in Appendix D.
@@
To evaluate whether models capture not only what individuals believe but also how their beliefs evolve, we operationalize reasoning through two proxy tasks: belief state inference (recovering stance and factor polarity from context) and belief dynamics update (predicting stance shifts and reweighting under new evidence). These tasks follow the tradition of modeling belief revision as a tractable proxy for underlying cognitive processes (gopnik2004theory; sloman2009causal). While other proxies could be envisioned, these two are the most direct operationalizations of individual reasoning trajectories, balancing interpretability and task difficulty. This motivates our benchmark’s two-task structure, detailed in Appendix D.
@@
Human data provide ecological validity: rich, idiosyncratic reasoning paths embedded in natural language. Synthetic agents, by contrast, provide controllability and scale: fully specified stance profiles and deterministic update rules allow stress-testing model adaptation under known ground truth. Together, the two tracks are complementary: humans as ecological baselines, synthetics as controlled stress tests. This mirrors dual-track designs in cognitive science and simulation benchmarks, where naturalistic and synthetic data jointly enhance validity and reproducibility (lake2017building; battaglia2018relational). Synthetic agents are not intended to replace human data but to serve as a complementary axis of evaluation. This dual-track design is what anchors HugAgent between ecological realism and controlled generalization tests.
@@
Instead, we adopt test–retest reliability as the human ceiling. Here, the same participant is re-sampled or re-interviewed, and the consistency of their own responses provides a direct measure of reliability. This practice is well established in psychology and survey research, and has been adopted in recent large-scale reasoning datasets facing similar challenges (park2023thousandpeople; toubia2025twin2k500datasetbuildingdigital). Compared to annotator baselines, test–retest reliability offers a more precise and ecologically valid upper bound for model performance, aligned with the benchmark’s goal of capturing intra-individual reasoning fidelity. This principle defines how we report the human ceiling in HugAgent.